OpenStudy (anonymous):

How fast must you drive to cover 700 miles in 13 hours? Round your answer to the nearest whole number.

OpenStudy (anonymous):

700miles/13hours=53.85miles/hour or approx 55mphin california

OpenStudy (anonymous):

700/13=54 mph

OpenStudy (anonymous):

@mark not 55

OpenStudy (anonymous):

i said in california maximux drivind lol

OpenStudy (anonymous):

distance=rate*time, so 700=r*13 or solving for rate =700miles/13hours.

OpenStudy (anonymous):

maximu drive in california is 55 mph,,lol

OpenStudy (anonymous):

so the correct ans is 53.85 or 54 approximately

OpenStudy (anonymous):

700/13=53.84 Answer-->54mph